OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line with slope 9/7 and containing the midpoint of the segment whose end points are (2, -3) and (-6, 5).

OpenStudy (lgbasallote):

go @apoorvk pwn it ;)

OpenStudy (lgbasallote):

we're all watching you...

OpenStudy (anonymous):

lol

OpenStudy (apoorvk):

Wow. Great. let's do this then. -_- O shell from the sea, with or without an oyster inside... Do you know how to find the mid-point of a segment whose two end points are given?

OpenStudy (anonymous):

nope... ._. vacation stole my brain

OpenStudy (apoorvk):

Oh no problem, buy a new one from Walmart then. (come cheap these days)

OpenStudy (apoorvk):

Lol jokes apart,

OpenStudy (anonymous):

ehh okay

OpenStudy (apoorvk):

The midpoint 'M' of segment between two points (x1,y1) and (x2,y2) is: \[M = (\frac{x_1 + x_2}2 , \frac{y_1 + y_2}2)\] Does this^ bring back some distant faded memories of a certain "mid-point formula"? Can you use this?

OpenStudy (anonymous):

no it does not bring memories

OpenStudy (apoorvk):

Well, forget the non-existent memories then, can you just this formula to find out the mid-point of the two points (2,-3) and (-6,5) ? You just have to substitute the values.

OpenStudy (anonymous):

kay

OpenStudy (apoorvk):

Hmm, so what would the midpoint be?

OpenStudy (anonymous):

2,-3

OpenStudy (apoorvk):

*facepalm* (2,-3) is one of those points, how can it be the mid-point as well??

OpenStudy (anonymous):

-2,1

OpenStudy (anonymous):

.-.

OpenStudy (apoorvk):

Ahaa... right now! Good good. Now the next part. The equation of a line in its slope-intercept form is: \[y = mx +c\] Here, m = slope, and c = constant. For, now you know the slope of the line required is 9/7. So, this lines equation in terms of the unknown 'c' would be: \[y = \frac 9 7 x + c\] Okay with this^?

OpenStudy (anonymous):

wait... there cant be no fractions

OpenStudy (apoorvk):

Yeah, no probs, we will rearrange the whole thing later on.. Just do as I say for now.

OpenStudy (anonymous):

ok

OpenStudy (apoorvk):

so, we need to find out 'c' now right?

OpenStudy (anonymous):

right

OpenStudy (apoorvk):

Okay so we know, that this line passes through that mid-point (-2,1) that we found out earlier. So, (-2,1) must satisfy the equation of the line. so let's substitute x=-2 and y=1 in the equation of the line we have so far, and find out 'c'. can you do this? i think you can...

OpenStudy (anonymous):

i cant..

OpenStudy (anonymous):

1 = 9/7 (-2) + c

OpenStudy (apoorvk):

yeah, just like @jollysailorbold says^ I just put in x = -2, and y = 1 in the equation, and I get: \[1 = \frac 9 7 (-2) + c\] now, can you solve this and find out the value of 'c'?

OpenStudy (anonymous):

12?

OpenStudy (apoorvk):

\[ c = 1 + 2 \times\frac 9 7\]\[or, c = 1 + \frac {18} 7\] so, what's 'c' then?

OpenStudy (anonymous):

1 18/7 ??

OpenStudy (apoorvk):

\[c = 1 + \frac {18} 7 = \frac{7 + 18}{7} = ??\]

OpenStudy (anonymous):

:/ i dont get itt

OpenStudy (apoorvk):

can you add two fractions? Do you know how this is done?

OpenStudy (anonymous):

yes..

OpenStudy (apoorvk):

So what's the trouble here? You just need to add 1 and 18/7 ! Let's go slow-motion then. \[c = 1 + \frac {18} 7\]\[ =\frac1 1 + \frac {18} 7\]\[ = \frac {7\times 1} {7 \times 1} + \frac {18} 7\]\[ = \frac 7 7 + \frac {18} 7\]\[= \frac{7 + 18}{7} = ??\]

OpenStudy (anonymous):

25/7

OpenStudy (apoorvk):

Finally. -_- yeah, so we found out what Mr. 'c' is. so, in the equation of the line we had had in terms of x and y, that is, this one: \[y = \frac 9 7 x + c\] I just replace 'c' with it's obtained value. What equation do I have?

OpenStudy (anonymous):

hahahaha omg you just made me laugh

OpenStudy (apoorvk):

Do I have \[y = \frac 9 7 x + \frac {25} 7\] ??

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

the answers can be in fraction form

OpenStudy (apoorvk):

You will make me cry soon though. -_- jus' kidding. xD Okay, so we are almooooost there. What I have is the equation of the line in slope intercept form. But we can't have the fractions. So what do I do? there's a '7' in the denominator. Ohkay, ohkay, Mr. '7', you're goin' up. -.- I multiply both sides by 7. What do I get then?

OpenStudy (apoorvk):

can be or "can't" be?

OpenStudy (anonymous):

wait....waht doi do?

OpenStudy (anonymous):

omg this is like... only one question!!!! and like.... its been 1 hr!!!!!

OpenStudy (apoorvk):

I have: \[y = \frac 9 7 x + \frac {25} 7\] I multiply both sides by '7' I get: \[7y = 9x + 25\] that's^ what you need. -_-

OpenStudy (anonymous):

-.- i cant do math

OpenStudy (anonymous):

63x+175?

OpenStudy (apoorvk):

As for the time, I have been waiting for you to reply for almost 25 minutes. And I can't help it either, if I just give out an answer, then I guess the fresh brain you just bought from Walmart will remain fresh all it's life. -.- WHAT DID YOU DO?? Why did you multiply it by '7' again, I just got rid of the fractions!

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

so u already gave me the answer?

OpenStudy (apoorvk):

@seashell you are not weak at maths, or can't do it, you just lack confidence and the belief that you can do it. The day you get rid of this fear, you will start loving maths. believe me ;)

OpenStudy (apoorvk):

yeah, that "7x = 9y + 25" is the answer. Now be honest, did you understand what we did here, or is your mind still in a tailspin?

OpenStudy (anonymous):

both

OpenStudy (apoorvk):

Ohkay, great! Atleast you understood! :D take a break to calm your mind. :P

OpenStudy (anonymous):

kay:D